OpenStudy (anonymous):

What is the solution to the following equation? 5(2x − 14) + 23 = 7x − 14 a.11 b.14 c.30 d.36

OpenStudy (anonymous):

A

OpenStudy (anonymous):

5(2x − 14) + 23 = 7x − 14 10x-70+23=7x-14 10x-7x=70-23-14 3x=33 X=11

OpenStudy (anonymous):

Thank you(:

OpenStudy (anonymous):

Thank you(: